Devise a strategy to derive an expression of Electric Field at a point R due to an Electric
Dipole by gausss law or columbs law
1
Expert's answer
2020-11-23T10:16:09-0500

Let the x component of the electric field strength at point P with Cartesian co-ordinates (x, y, 0)


"E_1=\frac{2p\cos\theta}{4\pi \epsilon_o r^3}" along PD


"E_2=\frac{p\sin\theta}{4\pi \epsilon_o r^3}" along PC

Net, electric field intensity

"E=\sqrt{E_1^2+E_2^2}"

"\Rightarrow E=\sqrt{(\frac{2p\cos\theta}{4\pi \epsilon_o r^3})^2+(\frac{p\sin\theta}{4\pi \epsilon_o r^3})^2}"

"\Rightarrow E =\frac{p}{4\pi \epsilon_o r^3} \sqrt{4\cos^2\theta +\sin^2\theta}"

"\Rightarrow E =\frac{p}{4\pi \epsilon_o r^3} \sqrt{3\cos^2\theta +1}"